Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#247 closed flight processing (fixed)

HY05/02, flight day 153/2009, Ayrshire Coalfields

Reported by: jatt Owned by: knpa
Priority: alpha 4 medium Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors: mark1

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/uk/HY05_02-2009_153_Ayrshire_coalfields

Data arrived from ARSF via SATA disk M on 9th June 2009.

Scientific objective: N/K

Priority: α4M

PI: C Fleming

Sensors:

  • Eagle
  • Hawk
  • Leica LIDAR (14/05/2020)
  • Photography (15/01/2010)

Attachments (1)

GB08_15rollVsTime.png (38.0 KB) - added by knpa 10 years ago.

Download all attachments as: .zip

Change History (39)

comment:1 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:2 Changed 10 years ago by knpa

There were three separate sorties on the flight day, each with a different logsheet. Am dividing the file structure (on just my scratch space for now) into 3 trees, one for each site.

comment:3 Changed 10 years ago by knpa

Applanix processing completed

comment:4 Changed 10 years ago by knpa

Both E+H appear to have SCT offsets which need to be corrected by hand for every flightline.

comment:5 Changed 10 years ago by knpa

Line 9 has an uncorrectable along track shift along it's south edge. Cause unknown.

comment:6 Changed 10 years ago by knpa

Line 1 has a stretch along it's interior, the line fell off the DEM to the north, am extending DEM.

comment:7 Changed 10 years ago by knpa

Extended DEM to the north with panels ns63 and ns73 and have eliminated stretch from Line 1.

comment:8 Changed 10 years ago by knpa

All Hawk flightlines have a regular drop of scan lines resulting in approx. 7 across-track black lines 10 pixels wide.

comment:9 Changed 10 years ago by knpa

Line 5 was causing azimport to seg fault. Error was determined to be caused by an incorrect value for the number of scan lines in the header file (given as 10). This has been changed to the correct number calculated from the size of the raw file (result = 7283 scan lines).

comment:10 Changed 10 years ago by knpa

#note# Above message is referring to Hawk Line 5.

comment:11 Changed 10 years ago by knpa

Eagle Line 14 - No Sync information - needs to be corrected manually

comment:12 Changed 10 years ago by mggr

Mail from Dr Ferrier:

My name is Dr graham ferrier, i am co-PI on HY05-02. Unfortunately i
wasn't able to be on the ground at the time of the overflight. I am
therefore trying to get up to the site relatively soon to take some
field spectral measurements of calibration targets. Would you have any
quicklook photographs of the dataset to enable me to target my field
campaign.

Kevin has been working on this so will collect some overview images.

comment:13 Changed 10 years ago by knpa

.png overview shots sent to PI

comment:14 Changed 10 years ago by knpa

SCT offsets for all flightlines for the three sites found (~/backup/sctHY05_02).

Awaiting arrival of Vectors before we can proceed to delivery checking.

comment:15 Changed 10 years ago by knpa

e-mailed PI to enquire whether or not they want to wait for vectors

comment:16 Changed 10 years ago by knpa

PI replied they are happy to wait for delivery upon final check.

Changed 10 years ago by knpa

comment:17 Changed 9 years ago by mark1

Vectors have arrived.

comment:18 Changed 9 years ago by knpa

Awaiting delivery check

comment:19 Changed 9 years ago by emca

  • Owner changed from knpa to emca
  • Status changed from assigned to new

comment:20 Changed 9 years ago by emca

Delivery checking complete.
Changes made to read_me and file names.
Ready to deliver.

comment:21 Changed 9 years ago by knpa

  • Owner changed from emca to knpa

comment:22 Changed 9 years ago by knpa

HYPERSPECTRAL DELIVERED TODAY

comment:23 Changed 9 years ago by knpa

  • Owner knpa deleted

comment:24 Changed 9 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:25 Changed 9 years ago by mark1

  • Description modified (diff)

comment:26 Changed 9 years ago by mark1

  • Description modified (diff)

Digital Photography dispatched to PI on 15 January 2010

comment:27 Changed 9 years ago by mark1

  • Other processors set to mark1

Started Lidar processing

comment:28 Changed 9 years ago by mark1

Delivery directories (LIDAR) created at:

~airborne/workspace/HY05_02-2009_153_Ayrshire_coalfields/HY05_02a-2009_153_Ayrshire_coalfields/delivery/20100427

~airborne/workspace/HY05_02-2009_153_Ayrshire_coalfields/HY05_02b-2009_153_Ayrshire_coalfields/delivery/20100428

~airborne/workspace/HY05_02-2009_153_Ayrshire_coalfields/HY05_02c-2009_153_Ayrshire_coalfields/delivery/20100428

comment:29 Changed 9 years ago by knpa

  • Owner changed from knpa to mark1

comment:30 Changed 9 years ago by emca

Delivery checked the three LIDAR deliveries - all data looks good.
Ready to deliver.

comment:31 Changed 9 years ago by mark1

  • Description modified (diff)

LIDAR data dispatched to PI on 14/05/2020

comment:32 Changed 9 years ago by mark1

Have moved the directory structure from:
HY05_02-2009_153_Ayrshire_coalfields/HY05_02*-2009_153_Ayrshire_coalfields
to
HY05_02-2009_153*_Ayrshire_coalfields

Have rsynced the lidar and photos over to ~arsf.

comment:33 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

Ready for archiving.

comment:34 Changed 9 years ago by knpa

Begining archiving.
Everything appears to be rsynced and the projects seems to be correctly split into it's three constituents, but will double check.

comment:35 Changed 9 years ago by knpa

  • Owner changed from mark1 to knpa

Transferred to, and backed up by, NEODC.

Have moved project from repository to archive disk.

comment:36 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace version and closing ticket.

comment:37 Changed 2 years ago by dac

Original delivery have incorrect wavelengths for Eagle. Data reprocessed with APL are OK.

comment:38 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.